OpenStudy (anonymous):

Verify the equation: 2. (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos^2 x

hartnn (hartnn):

1) distribute sin x 2) write tan x = sin x/ cos x, cot x = cos x/sin x.

OpenStudy (anonymous):

Can we start with the right side?

OpenStudy (anonymous):

Or no? Because I think I'm supposed to match right to left

hartnn (hartnn):

is it asked to march right to left ? that would be not as easy as to go from left to right...

OpenStudy (anonymous):

Yeah I know! haha but that's what it says lol

OpenStudy (sirm3d):

\[1-2\cos^2 x=1-\cos^2 x - \cos^2 x\\=\sin^2 x - \cos^2 x\\\]

hartnn (hartnn):

ok, then you can first use the formula, 1= \(\sin^2x + \cos^2x\)

OpenStudy (anonymous):

okay so sin^2 x+cos^2 x- 2 cos^2 x

hartnn (hartnn):

yes, which is sin^2x-cos^2x, isn't it ?

hartnn (hartnn):

now whats done is : \(\dfrac{\sin^2x \cos x}{\cos x} - \dfrac{\cos^2x \sin x}{\sin x} \)

hartnn (hartnn):

then 1)factor out sin x from numerator of both terms, 2) write sin x/cos x = tan x in 1st term and cos x/sin x =cot x in 2nd term. and you're done...

OpenStudy (anonymous):

okay so would I use the tanx=sinx/cosx and cotx=cosx/sinx formulas?

OpenStudy (anonymous):

oh you just wrote that haha

OpenStudy (anonymous):

okay let me try

hartnn (hartnn):

sure, take your time :)

OpenStudy (anonymous):

Would it be tan^2x cos x - cot^2 x sin x? I feel like that might be wrong

hartnn (hartnn):

tillwhich step are u sure, you got correct ?

OpenStudy (anonymous):

I'm not sure how to factor out sin x

hartnn (hartnn):

\(\dfrac{\sin^2x \cos x}{\cos x} - \dfrac{\cos^2x \sin x}{\sin x} \\ = \sin x [\dfrac{\sin x \cos x}{\cos x} - \dfrac{\cos^2x }{\sin x}] \\ =\sin x[\dfrac{\sin x }{\cos x} \times \cos x - \dfrac{\cos x }{\sin x} \times \cos x]\) now ?

OpenStudy (anonymous):

Oh! soo... sinx(tanx cosx- cot x cos x) You just made this whole verifying thing so much! Easier! I think that's where I'm messing up is factoring out! Thank you so much! :)

hartnn (hartnn):

welcome and keep practising ^_^